calculate the number of atoms in the following substances.

i) 13 g of zinc, Zn

ii) 5.6 g of nitrogen gas, N2

iii) 8.5g or ammonia gas, NH3

iv) 14.2 g of chlorine gas,Cl2


1
Expert's answer
2022-04-29T11:39:02-0400

1. M(Zn)=65 g/mol

n(Zn)=13/65=0.2 mol

So, there are 0.2x 6.02x1023=1.204 x1023 atoms.

2. M(N2)=14x2=28 g/mol

n(N2)5.6/28=0.2 mole

So, there are 0.2x 6.02x1023=1.204 x1023 molecules. 1 molecula contains 2 atoms. So there are 1.204x1023x2=2.408x1023 atoms.

3.M(NH3)=14+3x1=17 g/mol

n(NH3)=8.5/17=0.5 mole

So, there are 0.5x 6.02x1023=3.01x1023molecules. 1 molecula contains 4 atoms. So there are 3.01x1023x4=12.04x1023 atoms.

4.M(Cl2)=35.5x2=71 g/mol

n(Cl2)=14.2/71=0.2 mole. So, there are 0.2x 6.02x1023=1.204 x1023 molecules. 1 molecula contains 2 atoms. So there are 1.204x1023x2=2.408x1023 atoms.